0.73 g of hydrogen chloride gas is dissolved in 1000 cm3 of distilled water. Hydrogen chloride dissolves completely in water. What is the pH value of the resulting solution ?



1
Expert's answer
2021-08-18T03:34:01-0400

Molar mass of hydrogen chloride = 36.46

= 0.73/36.46

= 0.02 mol

Molarity = 0.02/1

= 0.02mol/L

Log (0.02) = -1.7

= 14 -1.7 = 12.3